OpenStudy (anonymous):

How would you factor x^2 + 2x - 3 ?

OpenStudy (ganpat):

x2 + 2x - 3 = 0 x2 + 3x -x -3 = 0 x (x + 3) - 1 (x + 3) = 0 (x - 1) (x + 3) = 0 x = 1 or x = -3